OpenStudy (anonymous):

I OMG my questions for math get harder and harder I need help solving this problem The demand for Sigma Mu Fraternity plastic brownie dishes is q(p) = 361,201 − (p + 2)2 where q represents the number of brownie dishes Sigma Mu can sell each month at a price of p¢. Use this function to determine the following. (1) the number of brownie dishes that Sigma Mu Fraternity can sell each month if the price is set at 90¢ (2) the number of brownie dishes they can unload each month if they give them away (3) the lowest price at which Sigma Mu will be unable to sell any dishes

OpenStudy (anonymous):

\[q(p) = 361,201 − (p + 2)^2 \] right?

OpenStudy (anonymous):

for number once, compute \[q(90) = 361,201 − (90 + 2)^2 \]

OpenStudy (anonymous):

for number two, compute \[q(0) = 361,201 − ( 2)^2 \]

OpenStudy (anonymous):

for number three, put \[361,201 − (p + 2)^2 =0\] and solve for \(p\)

OpenStudy (anonymous):

you get \[(p+2)^2=361,201\] \[p+2=\sqrt{361,201}=601\] \[p=599\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw, hope the first two are clear, because i did not work them out, but they should be easy enough, especially with a calculator